# arXiv:2603.00706v1[econ.GN] 28 Feb 2026

### v1

@@ 4.1 Experimental Setup
In each condition (SI, TI), the entrepreneur engages in bilateral negotiations with each investor present. Importantly, in TI, investor $i$ cannot see the offers (or any agreement) made between the entrepreneur and investor $j$. The negotiation format is semi-structured: players can make or accept offers specifying a share of the realized startup value (between 0 and 100%) that the investor will receive in exchange for their investment. Further, investments are all or nothing. That is, if the investor and entrepreneur reach an agreement, it is for the full amount of their endowment. No structure is placed on who may propose first, or on the order of proposals. Players do not see each other and may not exchange verbal messages during the negotiations. At the end of each round, the results of all negotiations are announced to all players in a dyad/triad. See Fig. EC1-EC2 for screenshots of the negotiation interface. In each period, the players in a negotiation dyad/triad have 90 seconds in the SI treatment or 180 seconds in the TI treatment to reach an agreement.We chose 90 seconds based on the bilateral bargaining times commonly used in the experimental literature (Isoni et al. 2022). We opted for a doubling of the bargaining time in the TI treatments, relative to SI. The rationale for this design choice is that the TI treatments included two bilateral negotiations, between the entrepreneur and investor $i$, as well as between the entrepreneur and investor $j$. If no agreement is reached with any investor, then all players receive their outside options. In the two investor cases, the entrepreneur can secure deals with zero, one or both investors. If negotiations succeed with only one investor, then the excluded investor receives their initial endowment, while the entrepreneur and investor who did reach an agreement are paid according to the terms of the agreement and the realized value of the startup. The complete set of experimental instructions and procedures is reproduced in Appx. EC.2.
@@ 4.3 Bargaining Process
To better understand Result 1 we examine the bargaining processes in each treatment (detailed analysis of the bargaining process is available in the supplementary document (https://researchbox.org/3335&PEER_REVIEW_passcode=XITWXN). We first show that the differences in bargaining outcomes between SI and TI conditions are primarily driven by initial offers rather than differences in concession behavior. When entrepreneurs make more generous initial offers (i.e., propose larger shares to investors), they significantly increase the chances of successfully reaching an agreement. Conversely, higher initial demands (asking for a larger share) by investors tend to lower agreement rates, significantly so in the SI condition. Moreover, initial positions strongly predict final outcomes: more generous initial offers by entrepreneurs, as well as higher initial demands by investors, lead to investors ultimately obtaining larger equity shares and vice versa for entrepreneurs. Final agreements tend to converge around the midpoint between the entrepreneur’s and investor’s initial proposals, suggesting that both parties make substantial concessions. This suggests a strategic trade-off faced by negotiators: setting generous initial terms facilitates reaching an agreement but reduces one’s own final payoff. The stronger presence of this trade-off in SI helps explain the observed differences in agreement rates between SI and TI. The analysis also suggests that entrepreneurs and investors adjust their opening strategies based on relative bargaining power, which drives the observed final outcomes in PoorEnt and RichEnt. Investors demand roughly six percentage points less when entrepreneurs have a positive outside option (RichEnt) than when entrepreneurs have no outside options (PoorEnt), indicating some responsiveness to the shift in bargaining power. However, these adjustments are minimal in the TI condition – less than one percentage point – suggesting that entrepreneurs are unable to leverage their improved outside option effectively when negotiating with two investors. This lack of strategic adjustment in TI, combined with the convergence of outcomes towards the midpoint of the negotiators’ initial positions, explains the difference in final outcomes and highlights the limits of the entrepreneur’s leverage in multi-party bargaining.
@@ 5.2 Bargaining Process and Fairness
To better understand Result 2, in the supplementary file (https://researchbox.org/3335&PEER_REVIEW_passcode=XITWXN) we examine the bargaining behaviors in each treatment and under each contract. As before, we see that the differences in final outcomes emerge primarily from the differences in opening offers. Entrepreneurs make lower first offers to investors under Preferred Stock contracts relative to Common Stock contracts (except in the SI-PoorEnt scenario). This behavior is consistent with theory, which predicts that entrepreneurs should be compensated for the additional risk they carry under Preferred Stock contracts. However, investors do not appear to respond to contracts in the same way. Indeed, their opening demands are virtually identical across contracts in the SI treatments ($p=0.500$ and $p=0.858$). Even more surprisingly, in the TI scenarios investors make significantly more aggressive opening demands under Preferred Stock than under Common Stock (increasing from approximately 44% to approximately 49% in both PoorEnt and RichEnt conditions, both $p<0.01$). The latter behavior contradicts theoretical expectations but may be plausibly explained by peer-induced fairness – each investor’s desire to outperform their peer investor (See Ho and Su 2009, Ho et al. 2014, for similar behaviors in ultimatum games and in supply chains). Without a peer investor in SI scenarios, this aggressive behavior diminishes. To better understand the drivers of these behaviors, we next examine the fairness beliefs of the negotiators.
@@ 6.1 Off-Equilibrium Belief Modeling
To explore a potentially more accurate modeling approach, we reconsider the second key assumption in our model: namely, that the agreement between the entrepreneur and investor $i$ is unaffected by whether the entrepreneur reaches agreement with investor $j$. This assumption gives the entrepreneur theoretical leverage in the TI treatment, since the entrepreneur can walk away from one negotiation without adversely affecting the other. Although common in the literature (Horn and Wolinsky 1988, Yürükoğlu 2022), this assumption may oversimplify actual bargaining dynamics. An alternative assumption about off-the-equilibrium-path beliefs in TI is that disagreement with one investor implies disagreement with both investors. Under this revised assumption, the entrepreneur’s bargaining leverage is diminished, as walking away from one negotiation necessarily terminates the other, reducing the entrepreneur’s ability to pit investors against each other. Incorporating this modified assumption yields “Revised model (ii)” in Table 5 (model details and proofs are available at https://researchbox.org/3335&PEER_REVIEW_passcode=XITWXN). As shown in the last row of the table, this specification achieves a further reduction in MSE (by 31.82%) and, importantly, yields more plausible estimates of bargaining power ($\hat{\theta}_{i}$).While we do not report estimates for Preferred Stock contracts, the same general patterns emerge, but the estimates of $\theta_{i}$ are higher, indicating more investor bargaining power, consistent with our previous results. The takeaway from this exercise is that the standard modeling approach of Nash-in-Nash bargaining with equal power and independent bilateral agreements may not adequately capture the complexities of multi-party entrepreneurial negotiations. How parties form beliefs about outcomes in alternative negotiation scenarios is an important consideration, particularly in settings where agreements are not independent. Our revised model that incorporates more realistic off-equilibrium beliefs offers a better fit to the experimental data and provides more plausible estimates of bargaining power.
@@ 8 Robustness Regressions
EC.2 Experimental Protocol and Instructions
EC.2.1 Experimental Protocol

### v2

# Kagan, Hyndman and Qi

@@ 4.1 Experimental Setup
In each condition (SI, TI), the entrepreneur engages in bilateral negotiations with each investor present. Importantly, in TI, investor $i$ cannot see the offers (or any agreement) made between the entrepreneur and investor $j$. The negotiation format is semi-structured: players can make or accept offers specifying a share of the realized startup value (between 0 and 100%) that the investor will receive in exchange for their investment. Further, investments are all or nothing. That is, if the investor and entrepreneur reach an agreement, it is for the full amount of their endowment. No structure is placed on who may propose first, or on the order of proposals. Players do not see each other and may not exchange verbal messages during the negotiations. At the end of each round, the results of all negotiations are announced to all players in a dyad/triad. See Fig. EC1-EC2 for screenshots of the negotiation interface. In each period, the players in a negotiation dyad/triad have 90 seconds in the SI treatment or 180 seconds in the TI treatment to reach an agreement.We chose 90 seconds based on the bilateral bargaining times commonly used in the experimental literature (Isoni et al. 2022). We opted for a doubling of the bargaining time in the TI treatments, relative to SI. The rationale for this design choice is that the TI treatments included two bilateral negotiations, between the entrepreneur and investor $i$, as well as between the entrepreneur and investor $j$. If no agreement is reached with any investor, then all players receive their outside options. In the two investor cases, the entrepreneur can secure deals with zero, one or both investors. If negotiations succeed with only one investor, then the excluded investor receives their initial endowment, while the entrepreneur and investor who did reach an agreement are paid according to the terms of the agreement and the realized value of the startup. The complete set of experimental instructions and procedures is reproduced in Appx. EC.3.
@@ 4.3 Bargaining Process
To better understand Result 1 we examine the bargaining processes in each treatment (detailed analysis of the bargaining process is available in Appendix EC.4. We first show that the differences in bargaining outcomes between SI and TI conditions are primarily driven by initial offers rather than differences in concession behavior. When entrepreneurs make more generous initial offers (i.e., propose larger shares to investors), they significantly increase the chances of successfully reaching an agreement. Conversely, higher initial demands (asking for a larger share) by investors tend to lower agreement rates, significantly so in the SI condition. Moreover, initial positions strongly predict final outcomes: more generous initial offers by entrepreneurs, as well as higher initial demands by investors, lead to investors ultimately obtaining larger equity shares and vice versa for entrepreneurs. Final agreements tend to converge around the midpoint between the entrepreneur’s and investor’s initial proposals, suggesting that both parties make substantial concessions. This suggests a strategic trade-off faced by negotiators: setting generous initial terms facilitates reaching an agreement but reduces one’s own final payoff. The stronger presence of this trade-off in SI helps explain the observed differences in agreement rates between SI and TI. The analysis also suggests that entrepreneurs and investors adjust their opening strategies based on relative bargaining power, which drives the observed final outcomes in PoorEnt and RichEnt. Investors demand roughly six percentage points less when entrepreneurs have a positive outside option (RichEnt) than when entrepreneurs have no outside options (PoorEnt), indicating some responsiveness to the shift in bargaining power. However, these adjustments are minimal in the TI condition – less than one percentage point – suggesting that entrepreneurs are unable to leverage their improved outside option effectively when negotiating with two investors. This lack of strategic adjustment in TI, combined with the convergence of outcomes towards the midpoint of the negotiators’ initial positions, explains the difference in final outcomes and highlights the limits of the entrepreneur’s leverage in multi-party bargaining.
@@ 5.2 Bargaining Process and Fairness
To better understand Result 2, in Appendix EC.4 we examine the bargaining behaviors in each treatment and under each contract. As before, we see that the differences in final outcomes emerge primarily from the differences in opening offers. Entrepreneurs make lower first offers to investors under Preferred Stock contracts relative to Common Stock contracts (except in the SI-PoorEnt scenario). This behavior is consistent with theory, which predicts that entrepreneurs should be compensated for the additional risk they carry under Preferred Stock contracts. However, investors do not appear to respond to contracts in the same way. Indeed, their opening demands are virtually identical across contracts in the SI treatments ($p=0.500$ and $p=0.858$). Even more surprisingly, in the TI scenarios investors make significantly more aggressive opening demands under Preferred Stock than under Common Stock (increasing from approximately 44% to approximately 49% in both PoorEnt and RichEnt conditions, both $p<0.01$). The latter behavior contradicts theoretical expectations but may be plausibly explained by peer-induced fairness – each investor’s desire to outperform their peer investor (See Ho and Su 2009, Ho et al. 2014, for similar behaviors in ultimatum games and in supply chains). Without a peer investor in SI scenarios, this aggressive behavior diminishes. To better understand the drivers of these behaviors, we next examine the fairness beliefs of the negotiators.
@@ 6.1 Off-Equilibrium Belief Modeling
To explore a potentially more accurate modeling approach, we reconsider the second key assumption in our model: namely, that the agreement between the entrepreneur and investor $i$ is unaffected by whether the entrepreneur reaches agreement with investor $j$. This assumption gives the entrepreneur theoretical leverage in the TI treatment, since the entrepreneur can walk away from one negotiation without adversely affecting the other. Although common in the literature (Horn and Wolinsky 1988, Yürükoğlu 2022), this assumption may oversimplify actual bargaining dynamics. An alternative assumption about off-the-equilibrium-path beliefs in TI is that disagreement with one investor implies disagreement with both investors. Under this revised assumption, the entrepreneur’s bargaining leverage is diminished, as walking away from one negotiation necessarily terminates the other, reducing the entrepreneur’s ability to pit investors against each other. Incorporating this modified assumption yields “Revised model (ii)” in Table 5 (model details and proofs are available in Appendix EC.2). As shown in the last row of the table, this specification achieves a further reduction in MSE (by 31.82%) and, importantly, yields more plausible estimates of bargaining power ($\hat{\theta}_{i}$).While we do not report estimates for Preferred Stock contracts, the same general patterns emerge, but the estimates of $\theta_{i}$ are higher, indicating more investor bargaining power, consistent with our previous results. The takeaway from this exercise is that the standard modeling approach of Nash-in-Nash bargaining with equal power and independent bilateral agreements may not adequately capture the complexities of multi-party entrepreneurial negotiations. How parties form beliefs about outcomes in alternative negotiation scenarios is an important consideration, particularly in settings where agreements are not independent. Our revised model that incorporates more realistic off-equilibrium beliefs offers a better fit to the experimental data and provides more plausible estimates of bargaining power.
@@ 8 Robustness Regressions
EC.2 Theory with Alternative Belief
We discuss the models under alternative belief about the disagreement points in this appendix.
EC.2.1 The Two Investors Model Under Common Stock Contract
In this section, we present the results and the proofs with the general bargaining powers of the investor(s) relative to the entrepreneur under the alternative disagreement point of the entrepreneur set as $d_{\text{e}}^{-i}=d_{\text{e}}$ under the Common Stock contract. All other model settings are the same as in Section  3.1. We note that the alternative specification of the entrepreneur’s disagreement point does not affect the result in the single investor bargaining problem. It only affects the two investor bargaining problem. The investments $I_{i}$ and the share $s_{i}$ maximize the following Nash product simultaneously:
$\max_{I_{i}\in[0,e],~s_{i}\in[0,1]}$~ $\left[\pi_{i}(\bm{I},\bm{s})-d_{i}\right]^{\theta_{i}}\left[\pi_{\text{e}}(\bm{I},\bm{s})-d_{\text{e}}^{-i}\right]^{1-\theta_{i}}$ (EC-66)
Proposition EC.2.1 (Two investors under alternative belief)
  • When ${\mu_{\alpha}}\geq\max\bigg\{\frac{3-2\theta_{1}-\theta_{1}\theta_{2}}{2(1-\theta_{1})}-\frac{d_{\text{e}}(\theta_{1}-\theta_{1}\theta_{2})}{e(1-\theta_{1})},~\frac{3-2\theta_{2}-\theta_{1}\theta_{2}}{2(1-\theta_{2})}-\frac{d_{\text{e}}(\theta_{2}-\theta_{1}\theta_{2})}{e(1-\theta_{2})}\bigg\}$, there exists an equilibrium bargaining outcome in which both investors invest with $I_{i}^{TI}=\nicefrac{{e}}{{2}}$ for $i\in\{1,2\}$, and the equilibrium share of investor $i$ is
    $s_{i}^{TI}=\frac{1-2\theta_{i}+2{\mu_{\alpha}}\theta_{i}+(1-2{\mu_{\alpha}})\theta_{1}\theta_{2}}{2{\mu_{\alpha}}(1-\theta_{1}\theta_{2})}-\frac{d_{\text{e}}(\theta_{i}-\theta_{1}\theta_{2})}{e{\mu_{\alpha}}(1-\theta_{1}\theta_{2})}.$ (EC-67)
  • There exists an equilibrium bargaining outcome in which only Investor $i$ invests when ${\mu_{\alpha}}<\frac{2-\theta_{i}}{1-\theta_{i}}-\frac{2d_{\text{e}}\theta_{i}}{e(1-\theta_{i})}$. The equilibrium investment level $I_{i}^{TI}=e$ and $I_{j}^{TI}=0$ for $i,j\in\{1,2\}$ and $i\neq j$, and the equilibrium share of Investor $i$ is
Proof of Proposition EC.2.1. Recall that the expected profit of the entrepreneur is
$\pi_{\text{e}}(\bm{I},\bm{s})={\mu_{\alpha}}(I_{1}+I_{2})(1-s_{1}-s_{2}),$ (EC-68)
and the expected profit of Investor $i$ is
$\pi_{i}(\bm{I},\bm{s})={\mu_{\alpha}}(I_{1}+I_{2})s_{i}+\frac{e}{2}-I_{i}.$ (EC-69)
The disagreement point of the entrepreneur when negotiating with Investor $i$ is $d_{\text{e}}^{-i}=d_{\text{e}},$ which the value of the outside option when the entrepreneur does not reach agreement with either investor. The disagreement point of Investor $i$ is $d_{i}=\nicefrac{{e}}{{2}}$ since the investor has $\nicefrac{{e}}{{2}}$ units of capital as the endowment. We first solve the bargaining problem between the entrepreneur and Investor 11 as specified in (EC-66). Following the similar analysis as in the proof of Proposition EC.1.1, we have that
$\pi_{1}(\bm{I},\bm{s})-d_{1}=\theta_{1}\left(\pi_{1}(\bm{I},\bm{s})+\pi_{\text{e}}(\bm{I},\bm{s})-d_{1}-d_{\text{e}}^{-1}\right);$ (EC-70)
Note that the best-response investment level
[…]
By Eq. (EC-70), the best-response share for Investor $1$ is
[…]
Solving the system of the best-response functions Eqs. (EC-73) through (EC-82), we have that if ${\mu_{\alpha}}\geq\max\bigg\{\frac{3-2\theta_{1}-\theta_{1}\theta_{2}}{2(1-\theta_{1})}-\frac{d_{\text{e}}(\theta_{1}-\theta_{1}\theta_{2})}{e(1-\theta_{1})},~\frac{3-2\theta_{2}-\theta_{1}\theta_{2}}{2(1-\theta_{2})}-\frac{d_{\text{e}}(\theta_{2}-\theta_{1}\theta_{2})}{e(1-\theta_{2})}\bigg\}$, there exists an equilibrium in which both investors invest $I_{i}^{TI}=e$ with the share for Investor $i$ as
$s_{i}^{TI}=\frac{1-2\theta_{i}+2{\mu_{\alpha}}\theta_{i}+(1-2{\mu_{\alpha}})\theta_{1}\theta_{2}}{2{\mu_{\alpha}}(1-\theta_{1}\theta_{2})}-\frac{d_{\text{e}}(\theta_{i}-\theta_{1}\theta_{2})}{e{\mu_{\alpha}}(1-\theta_{1}\theta_{2})}.$
$\blacksquare$
EC.2.2 The Two Investors Model Under Preferred Stock Contract
In this section, we present the results and the proofs with the general bargaining powers of the investor(s) relative to the entrepreneur under the alternative disagreement point of the entrepreneur set as $d_{\text{e}}^{-i}=d_{\text{e}}$ under the Preferred Stock contract. All other model settings are the same as in Section 5. We note that the alternative specification of the entrepreneur’s disagreement point does not affect the result in the single investor bargaining problem. It only affects the two investor bargaining problem. Similar to the analysis before, we can show that the equilibrium bargaining share will not be too extreme such that in any equilibrium bargaining outcome $(\bm{I},\bm{s})$, the following conditions are satisfied: $\alpha_{H}(I_{1}+I_{2})(1-s_{i})\geq I_{3-i},~i=1,2.$ Therefore, the expected profit of the entrepreneur is
$\pi_{\text{e}}(\bm{I},\bm{s})$ $=\alpha_{H}(I_{1}+I_{2})p-\sum_{i=1}^{2}\max\bigg\{\alpha_{H}(I_{1}+I_{2})s_{i},I_{i}\bigg\}p,$ (EC-83)
and the expected profit of Investor $i$ as
$\pi_{i}(\bm{I},\bm{s})$ $=\max\bigg\{\alpha_{H}(I_{1}+I_{2})s_{i},I_{i}\bigg\}p+e-I_{i}p.$ (EC-84)
The (alternative) disagreement point of the entrepreneur when negotiating with Investor $i$ is $d_{\text{e}}^{-i}=d_{\text{e}}$, and the disagreement point of Investor $i$ is $d_{i}=\nicefrac{{e}}{{2}}$ since the investor has $\nicefrac{{e}}{{2}}$ units of capital as the endowment. Then, the investments $\bm{I}$ and the shares $\bm{s}$ maximize the following Nash product simultaneously:
$\max_{I_{i}\in[0,e],~s_{i}\in[0,1]}$~ $\left[\pi_{i}(\bm{I},\bm{s})-d_{i}\right]^{\theta_{i}}\left[\pi_{\text{e}}(\bm{I},\bm{s})-d_{\text{e}}^{-i}\right]^{1-\theta_{i}}$ (EC-85)
Proposition EC.2.2 (Two investors under alternative belief)
  • There exists an equilibrium bargaining outcome in which both investors invest; i.e., $\tilde{I}_{i}^{TI}=\nicefrac{{e}}{{2}}$ for $i\in\{1,2\}$, and the equilibrium share of investor $i$ is as follows.
    • If $\alpha_{H}\geq\max\left\{\frac{3}{2}+\frac{(1-\theta_{1})\theta_{2}}{2(1-\theta_{2})}-\frac{d_{\text{e}}(1-\theta_{1})\theta_{2}}{ep(1-\theta_{2})},\frac{3}{2}+\frac{(1-\theta_{2})\theta_{1}}{2(1-\theta_{1})}-\frac{d_{\text{e}}(1-\theta_{2})\theta_{1}}{ep(1-\theta_{1})}\right\}$,
      $\tilde{s}_{1}^{TI}=\frac{\theta_{1}(2\alpha_{H}(1-\theta_{2})+\theta_{2}-2)+1}{2\alpha_{H}(1-\theta_{1}\theta_{2})}-\frac{d_{\text{e}}\theta_{1}(1-\theta_{2})}{\alpha_{H}ep(1-\theta_{1}\theta_{2})},$ (EC-86)
      $\tilde{s}_{2}^{TI}=\frac{\theta_{2}(2\alpha_{H}(1-\theta_{1})+\theta_{1}-2)+1}{2\alpha_{H}(1-\theta_{1}\theta_{2})}-\frac{d_{\text{e}}(1-\theta_{1})\theta_{2}}{\alpha_{H}ep(1-\theta_{1}\theta_{2})}$ (EC-87)
Proof of Proposition EC.2.2. We solve the bargaining problem between the entrepreneur and Investor 11 as specified in (EC-85) for the best-response investment level and the share of the investor. We first note that if $s_{1}<\frac{I_{1}}{\alpha_{H}(I_{1}+I_{2})}$, it follows that $\pi_{1}(\bm{I},\bm{s})=\nicefrac{{e}}{{2}}$ and the Nash product for the bargaining between the entrepreneur and Investor $1$ in problem (EC-85) is zero. In the following analysis, we restrict attention to the case where Investor $1$’s share $s_{1}\geq\frac{I_{1}}{\alpha_{H}(I_{1}+I_{2})}$ and later verify that the equilibrium bargaining outcome leads to a strictly positive Nash product. In this case, by the first order condition, we have that
$\pi_{1}(\bm{I},\bm{s})-d_{1}=\theta_{1}\left(\pi_{1}(\bm{I},\bm{s})+\pi_{\text{e}}(\bm{I},\bm{s})-d_{1}-d_{\text{e}}^{-1}\right);$ (EC-88)
Note that the best-response investment level
[…]
By Eq. (EC-88), the best-response share for Investor $1$ is
[…]
Solving the system of the best-response functions Eqs. (EC-91) through (EC-100), we have that there exists an equilibrium in which both investors invest $\tilde{I}_{i}^{TI}=\nicefrac{{e}}{{2}}$ with the share for Investor $i$ as follows.
  • If $\alpha_{H}\geq\max\left\{\frac{3}{2}+\frac{(1-\theta_{1})\theta_{2}}{2(1-\theta_{2})}-\frac{d_{\text{e}}(1-\theta_{1})\theta_{2}}{ep(1-\theta_{2})},\frac{3}{2}+\frac{(1-\theta_{2})\theta_{1}}{2(1-\theta_{1})}-\frac{d_{\text{e}}(1-\theta_{2})\theta_{1}}{ep(1-\theta_{1})}\right\}$,
    $\tilde{s}_{1}^{TI}=\frac{\theta_{1}(2\alpha_{H}(1-\theta_{2})+\theta_{2}-2)+1}{2\alpha_{H}(1-\theta_{1}\theta_{2})}-\frac{d_{\text{e}}\theta_{1}(1-\theta_{2})}{\alpha_{H}ep(1-\theta_{1}\theta_{2})},$
    $\tilde{s}_{2}^{TI}=\frac{\theta_{2}(2\alpha_{H}(1-\theta_{1})+\theta_{1}-2)+1}{2\alpha_{H}(1-\theta_{1}\theta_{2})}-\frac{d_{\text{e}}(1-\theta_{1})\theta_{2}}{\alpha_{H}ep(1-\theta_{1}\theta_{2})}$
$\blacksquare$
EC.3 Experimental Protocol and Instructions
EC.3.1 Experimental Protocol
@@
EC.4 Insights From the Bargaining Process
In this section, we delve into the bargaining process in order to better understand the drivers of our results in §4-5. We begin by reproducing the standard result in the bargaining literature that first offers typically have an anchoring effect on negotiated outcomes. Building on this result, we then show that there are noticeable treatment differences in opening offers, which, combined with anchoring, go a long way to explaining our key findings.
EC.4.1 Anchoring
Table EC1 reports the results of a series of random effects regressions analyzing how first offers (of both investors and entrepreneurs) affect the likelihood of successfully reaching an agreement and also on the investors’ shares conditional on an agreement. Here we focus on SI and TI treatment conditions, in which bargaining is one-dimensional (negotiators only negotiate equity percentage, and not the investment size). Note that because we focus on bilateral negotiations between individual investor $i$ and the entrepreneur we use investor $i$ share (as opposed to entrepreneur share) as our measurement.Table EC1 suggests strong anchoring effects for both agreements and shares. Higher opening offers to investors by entrepreneurs are significantly positively associated with agreement, while higher opening demands by investors are negatively (and, for SI, significantly) associated with agreements. That is, the more generous the entrepreneur, the more likely is an agreement, while the more demanding the investor, the less likely is an agreement. Turning now to the agreed investor shares, we see that both the offers made to and the demands made by investors are significantly positively associated with the investor’s final share. These regression results show that players generally face a trade-off in bargaining. The more generous they are to their bargaining partner, the more likely are they to reach an agreement, but the less favorable the agreement will be to themselves.
Table EC1: Anchoring Effects of Initial Offers on Agreements and Investor Shares SI           TI Dep. Var.: Agreement Reached Investor $i$’s Share        Agreement Reached        Investor $i$’s Share Initial Offer to Investor $i$ by Entrepreneur 0.005^*** (0.001) 0.291^*** (0.047) 0.007^*** (0.002) 0.399^*** (0.057) Initial Demand by Investor $i$ -0.006^*** (0.001) 0.384^*** (0.051) -0.001 (0.001) 0.129^*** (0.046) Constant 1.076^*** (0.051) 17.894^*** (3.785) 0.767^*** (0.106) 17.889^*** (3.340) $R^{2}$ 0.138 0.421 0.048 0.390 $N$ 1342 1104 838 724 Note: Random effects regression coefficients are reported (Standard errors in parentheses). All regressions include controls for treatment variables and standard errors are corrected for clustering at the session level. The number of observations ($N$) corresponds to the number of interactions in which both negotiators (entrepreneur and investor $i$) made at least one offer. * p<<0.10, ** p<<0.05, *** p<<0.01.
Result B.1
There are strong anchoring effects: first offers predict both the likelihood of agreement and the final share received conditional on an agreement. Negotiators face a trade-off. A more aggressive opening offer increases the chance of disagreement, but, conditional on an agreement, leads to a more favorable outcome.
EC.4.2 Bargaining Process: Initial Offers in PoorEnt and RichEnt
We next examine the initial offers made by investors and entrepreneurs in PoorEnt and RichEnt. In PoorEnt offers were one-dimensional and consisted of the share offered to the investor in exchange for an investment of 200 (in SI) or 100 (in TI). In Table EC2 we report summary statistics on the proposed share to the investor.
Table EC2: Opening Offers: Proposed Share to Investor for First Offer
Treatment condition Proposed Investment Entrepreneur’s Proposal Investor $i$’s Proposal
PoorEnt RichEnt PoorEnt vs. RichEnt $p$-value PoorEnt RichEnt PoorEnt vs. RichEnt $p$-value
SI 200 36.97 40.30 0.348 70.77 65.13 0.313
TI 100 23.78 23.31 0.807 44.29 44.54 0.989
Note: $p$-values are obtained from $t$-tests on the subject average of first offers for each investment amount.
On average, entrepreneurs offer investors 36.97% of the pie in exchange for an investment of 200 in SI and 23.78% in exchange for an investment of 100 TI; thus, while the funding amount is halved, the share offered to each investor drops by less than half as we go from SI to TI. A similar pattern emerges if we look at investors’ first offers. Further, the final share obtained by investor $i$ is 46.70% in SI PoorEnt and 32.60% in TI PoorEnt (51.65% in SI RichEnt and 32.82% in TI RichEnt, see Tables 2-3), which is close to the midpoint between the entrepreneur’s and investor’s opening offers, suggesting that both parties make substantial concessions to get to an agreement. These comparisons suggest that the main results in §4.2 are well explained by the investors’ and entrepreneurs’ initial negotiation strategies (opening offers), and not by the differences in bargaining dynamics.
Result B.2
Opening offers of both parties largely explain the differences in final outcomes in SI and TI. Poor entrepreneurs prefer larger investments, while rich entrepreneurs prefer smaller investments. Investors prefer smaller investments. Investors bargain more aggressively when the entrepreneur is rich.
EC.4.3 Differences Across Contracts
We conclude this section by examining the bargaining dynamics that cause the differences between Common vs. Preferred Stock contracts. Table EC3 shows summary statistics on first offers for each role and for each of our treatment conditions in the SI/TI treatment arm. There are several interesting observations. First, consistent with theory, entrepreneurs’ first offers to investors were generally lower under Preferred Stock contracts than under Common Stock contracts ($p<0.05$ in three out of four comparisons). The only scenario in which entrepreneurs do not increase their demands with Preferred (relative to Common) Stock contracts is SI-PoorEnt. This is consistent with Result 3 in §5, which also reports that Preferred Stock contracts disadvantage entrepreneurs in the PoorEnt treatment arms, with the effect being particularly stark in the SI condition.
Table EC3: First Offers (Share to Investor) by Bargaining Environment and Player Role
Entrepreneur’s First Offer Investor $i$’s First Offer
Common Preferred $p$-value Common Preferred $p$-value
PoorEnt 36.97 39.70 0.303 70.77 71.60 0.858
RichEnt 40.30 34.25 $\ll 0.01$ 65.13 66.71 0.500
(a) SI
Entrepreneur’s First Offer Investor $i$’s First Offer
Common Preferred $p$-value Common Preferred $p$-value
PoorEnt 23.78 20.48 $\ll 0.01$ 44.29 49.59 0.004
RichEnt 23.31 20.30 0.031 44.54 49.38 $\ll 0.01$
(b) TI
Note: $p$-values are derived from random effects regressions of first offers on treatment indicators, with standard errors corrected for clustering at the session level.
Second, investors’ opening demands were either approximately the same (SI) or significantly higher (TI) for Preferred relative to Common Stock contracts. In particular, in TI, investors increased their demands from an average of 44.29% to 49.59% in PoorEnt ($p=0.004$) and from 44.54 to 49.38% ($p\ll 0.01$) in RichEnt. That is, despite their reduced theoretical bargaining power with Preferred Stock contracts, investors adopted more extreme opening positions in the two investor (TI) setting. While we cannot provide definitive evidence for the driver of this behavior, it is consistent with peer-induced fairness (Ho and Su 2009, Ho et al. 2014) – each investor’s desire to outcompete the other investor, resulting in more aggressive bargaining positions. Conversely, the absence of a peer investor in the SI scenarios may lead to less aggressive investor behaviors.
Result B.3
Insufficient reflection of contract type in the share allocation (Result 2 in §5) is explained by insufficient adjustment of entrepreneurs’ opening offers in SI PoorEnt scenario, and by more aggressive opening offers by investors, particularly in TI scenarios.